# Aging-aware Energy Management for Residential Multi-Carrier Energy Systems

### v2

@@ 5.1 Case Study I: Day-ahead planning
For the EV timeseries comparison, Fig.8, the user’s mobility requirement leads to similar timeseries for all the planners. In general CPBDeg delays the charging to extend the V2G as much as possible, to avoid $\overline{SoC}_{\textrm{EV}}$. V2G mode is less frequent in the BNoDeg/CEmpDeg than in the CPBDeg because of this, being valid in both standard seasons. Finally, for the TESS, Fig. 9 shows that in the summer the CPBDeg has the smallest ST curtailment, because the chosen trajectory is the lowest of the three. BNoDeg and CEmpDeg chose almost the same trajectory for $SoC_{\textrm{TESS},t}$.
@@
The performance of each planner is summarized in Table 1. Starting with the grid cost $C_{\textrm{grid}}$, the best performer is the CEmpDeg in the summer and BNoDeg in the winter. The differences between first and second are less than 1 €. The worst performer is CPBDeg. For the total capacity fade $Q_{\textrm{loss}}$, CEmpDeg has the highest degradation in the summer and BNoDeg has the highest in the winter. The proposed aging aware CPBDeg achieves the lowest degradation in both seasons maintaining a reasonable $C_{\textrm{grid}}$.
@@
In general the planner that has the lowest $C_{\textrm{grid}}$ is the one with the highest $Q_{\textrm{loss}}$, but this is not always BNoDeg. The CEmpDeg fails to minimize the total capacity fade because of its model bias ($S^{M}_{a,t}\neq\tilde{S}^{M}_{a,t}$), in which it is rewarded to cycle at $\overline{SoC}_{b}$ and a calendar ageing independent of the $SoC_{b,t}$. Additionally the linear BNoDeg planner fails to fulfill its predictions because of the high number of rejected actions during summer. This is because when the BNoDeg plans for high-power discharges its predicted $\tilde{SoC}_{b,t}$ deviates from the true $SoC_{b,t}$ of the simulator $S^{M}_{a,t}$, wrongly depleting the ESS early (lower bound $\underline{SoC}_{b,t}$). The linear BNoDeg is the planner with the highest percentage of rejects, roughly 45% of the total time in the summer.
@@
Figure 10: Full equivalent cycles $FEC$ over time (left) and relative capacity fade $Q_{\textrm{loss}}$ over $FEC$ (right).
@@
However, when looking at the total storage usage, one must also analyze the number of cycles done by the $b$. Figure 10 presents the full equivalent cycles $FEC$ over time $t$, showing that the CPBDeg planners increase the EV usage $FEC_{\textrm{EV}}$ at the expense of doing less cycles with the BESS. When analyzing the capacity fade $Q_{\textrm{loss}}$ against the full eq. cycles $FEC$, Fig. 10, it is clear that the relative degradation per cycle ($\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$) of the CPBDeg is smaller in the winter than the used benchmarks. In summer, the BNoDeg has the lowest degradation per cycle due to its high number of rejects. The use of the EV s is the same for all planners, because the trajectory is mostly driven by user mobility. Finally, in both seasons CEmpDeg presents the highest $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$. This appears to be as a risky strategy due to a lack of consistency across season and objectives (minimizing degradation or minimizing grid costs). Lastly eventhough the capacity fade is not significant in $T=1$ month, daily optimization can have significant impact in the long term, as it was shown in [40, 39, 35]. As a final note, if the C-rate is increased ($\geq$1) and battery temperature $T$ is not constant the degradation on a daily basis can be significant.
Finally, the computational time for the different strategies is presented in Fig. 11. Each sample is the total computational time it takes to solve Algorithm 2. Unexpectedly, BNoDeg has the lowest and most consistent $t_{\textrm{comp}}$ distribution, i.e. the smallest standard deviation. However both CEmpDeg and CPBDeg planners have similar empirical distributions, maintaining overall fast computational time. Thus the increase in modeling accuracy of PBROM is not prohibitively expensive when compared to its empirical counterpart. This is to be expected as the empirical ageing model is also non-linear and non-convex.
@@
5.2 Case Study II: Managing different cathodes
To demonstrate the PB models’ flexibility and extended capabilities, the CPBDeg scheduler is tested using two similar battery packs of the same rated capacity $Q$ but using different cells. One is formed with NMC cells and the other with LFP. Since LFP cells have a lower rated capacity of $Q_{n}=2.3$ Ah and a lower $OCV$, the battery packs have more $N_{s/p}$ to have roughly the same pack-rated capacity as its NMC counterparts. The power limits $\overline{P}_{b,t},\ \underline{P}_{b,t}$ are also maintained to make an even comparison.
Figure 12: Cathode comparison simulation $SoC_{\textrm{EV}}$.
The PB aging models are suitable for both because they have graphite anodes [26, 15]. Nevertheless, they have different electrolytes. This is addressed by changing the electrolyte parameters in the model. Thus we use the same model equations but with different parameter values. This is a great advantage compared to the empirical fits presented in the literature. In the latter, the derived models are prone to overfitting to training conditions, delivering complex non-linear equations that can only be applied to specific chemistries, battery pack designs, and operating conditions. For our CPBDeg the physical parameters of the LFP cell used are taken from [27] and can be found in A, Tables 4 and 5.
@@ 5.2 Case Study II: Managing different cathodes
The simulation results are presented in Figs. 12 - 14. Starting with the EV, Fig. 12, the operation is similar except for a few days in summer and winter in which the LFP decides to have deeper discharges than its NMC counterpart. Moving forward to the BESS, Fig. 13, in summer the LFP scheduler is too ambitious generating early storage depletion roughly 40% of the time, 10% more than the NMC cells. During the winter total rejections are reduced to 25% and 10% for LFP and NMC respectively.
@@
Figure 14: Case Study 2 - LFP and NMC cells degradation analysis.
@@
Table 2 presents the summary of performance for both cell types. Overall the CPBDeg-LFP achieve lower totals costs and ageing despite their high number of rejected actions. During winter $C_{\textrm{grid}}$ stays the same with 37% less degradation. However, in the summer, just by changing the cell type from NMC to LFP $C_{\textrm{grid}}$ is reduced by 25%, similar to BNoDeg in Section 5.1, and roughly 25% less capacity fade. This shows that CPBDeg rightly exploits its physical information of the system to achieve better performance.
5.3 Case Study III: Managing aged and fresh batteries
To demonstrate the flexibility and extended capabilities of the CPBDeg planner the scheduler is tested using two battery packs: the first is the fresh battery pack of NMC cells of Section 5.1 and the second is the same pack but with cells aged at $SoH=90\%$. Only one benchmark is used: a BNoDeg with no $SoH$ update. Thus, the BNoDeg EMS sees a perfectly healthy cell with rated capacity when in reality the battery pack is aged 10%. The update is based on a 5% increase in series resistance $R_{b,0}$ and 10% decrease of the available Li content $z_{100\%,b}$ and its propagation with the equations of Section 3.1.2.
Figure 15 presents the degradation patterns for the 2 planners and different $SoH$. In the aged battery, the share of calendar ageing (against the total) is much smaller and thus the percentual $Q_{\textrm{loss}}$ is almost 35% smaller than in the new battery packs. In both seasons, the EV s patterns are similar for both BNoDeg and CPBDeg. The change in EV trajectories between aged and fresh cells shows slight increase $FEC$, due to smaller rated capacity, and a decrease in relative ageing, due to reduced calendar ageing. The impact on the BESS is more pronounced. In winter, CPBDeg has smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ (upper-right hand side of the graph) for both aged and new cells with BNoDeg doing more eq. cycles. In the winter CPBDeg still has a smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ in the new cell but not in the used cell, where this is achieved by the benchmark controller.
For the current simulation time $T$ of 1 month and average C-rate below 1C, the degradation slope may appear to be linear to the naked eye. However, in Fig 15 the trajectory of the aged battery is also presented and its average slope is smaller than the new battery degradation slope ($\frac{\Delta Q_{\textrm{loss},b}}{\Delta FEC_{b}}|_{SoH=90\%}<\frac{\Delta Q_{\textrm{loss},b}}{\Delta FEC_{b}}|_{SoH=100\%}$). This is due to the nonlinear dependency of $i_{\text{SEI},b,t}$ with $\sqrt{t}$. Thus it is clear that even though each trajectory might appear as linear, over longer simulation times of several months and years the dependency is non-linear.
Finally, when adding the costs to the analysis, summarized in Table 3, it is noticed that:
  • 1.
    CPBDeg achieves lower capacity fade $Q_{\textrm{loss}}$ than BNoDeg across all seasons and $SoH$.
  • 2.
    CPBDeg improves its total grid cost $C_{\textrm{grid}}$ as the cells degrade, in particular during summer.
  • 3.
    BNoDeg worsens its performance as the cells degrade. With higher $C_{\textrm{grid}}$, $Q_{\textrm{loss}}$ and model bias.
In summary, the performance of the proposed CPBDeg planner comparatively improves when using used cells, with respect to its linear BNoDeg benchmark.
@@ 5.3 Case Study III: Managing aged and fresh batteries
5.4 Limitations of the presented approach.
@@
    Even though the terminal set is more flexible than the standard literature the design is still arbitrary. It would be interesting to design the terminal conditions as part of a wider control scheme, such as seasonal optimization.
@@
    State observers are also necessary for all devices in an experimental application, to correctly feed back the states $S_{a,t}$ to the optimization.
@@
    The thermal flow-based models used in this paper assume a fixed building temperature decided in an external system. This is the assumption behind $P_{\textrm{load}}^{\textrm{th}}$.
@@
    Uncertainty handling is limited since re-optimization frequency is only 24hs. Future works include improving this through dynamic programming approaches and/or model predictive control.
  • 7.
    The local controls and protection of devices are also out of the scope. Their inclusion could lead to fault-triggered optimizations.
@@ 6 Conclusions & Discussion
In summary, this paper presents an optimization-based day-ahead planner for residential multi-carrier energy systems that uses PBROM models to integrate battery ageing. The proposed CPBDeg planner can handle different cathode chemistries as well as batteries in different ageing states. The planner does this with lower degradation than the benchmarks. This comes the expense of slightly higher computational times and grid cost.
@@
In the first case study it is shown how advanced PBROM can be used to reduce battery ageing while maintaining a reasonable grid cost, in accordance to the literature for standalone utility-scale applications [59, 24, 60, 35, 34]. This is because the planner can control the degradation per full equivalent cycle better than the benchmarks.
@@
In the second case study, the proposed planner is equipped with battery packs of different cathode chemistries and its performance is compared. This is out of the scope of most empirical degradation models. The LFP battery pack has a lower total $Q_{\textrm{loss}}$ as per established knowledge and can achieve lower grid costs $C_{\textrm{grid}}$ than its NMC counterpart. Even more so, when considering that the relative cost $c_{\textrm{loss}}$ used for both packs was the same, when in reality LFP packs have a lower cost than NMC packs. Thus the Section 5.2 is a conservative approximation and LFP packs have the potential to enable even lower operation costs. This is achieved just by changing some simple parameter models taken from the literature. This is an important feature because it allows the EMS to fully exploit any battery pack at hand. During the last case study it was shown how the proposed EMA handles aged and new batteries seamlessly, even improving performance. Its linear counterpart however is unaware of the degradation, increasing the number of rejected actions. Summing up reducing the grid costs can be achieved by using LFP battery packs or modifying objective weights, the latter at the expense of higher degradation.
It is worth noting again that in this work only one optimization is done per day for all models (benchmarks and novel). If re-optimizations occur frequently as in MPC, model bias can be mitigated. Future works will seek to implement the proposed models in an economic MPC approach. Moreover, for physical setups the proposed physics-based approach requires a non-linear observer to identify internal states and a system identification algorithm to parametrize the models. Adaptive control techniques/ online learning techniques are crucial for scaling implementation. For the case of the BESS initial tests and model identification can be done offline before start-up, and even offline during operation if historic data is continuously stored. However, for EV s models, parametrization deems a challenge since a previously unknown car may appear or due to unknown driving conditions and profiles. Thus, the development of effective and accurate observers to identify and parametrize PBROM online automatically is crucial. This allows the EMS predictions to be closer to reality, minimizing setpoint rejection by the BMS [17, 18, 48].
Another limitation of the proposed EMA is the modelling of the thermal carrier. In this paper, all thermal devices are assumed linear, and the horizon window $H$ is only 48hs. Thus seasonal and monthly variations can not be handled properly [42] as in other studies, like long-term hydrothermal dispatch problems [61]. Future works will address this shortcoming and improve the thermal models for seasonal planning [20, 30]. Finally, it would be interesting to investigate the dynamic linearization of PB models during operation to simplify the OCP for certain applications, potentially improving computational time and scalability.

### v3

@@ 5.1 Case Study I: Day-ahead planning
The performance of each planner is summarized in Table 1. Starting with the grid cost $C_{\textrm{grid}}$, the best performer is the CEmpDeg in the summer and BNoDeg in the winter. The differences between the first and second are less than 1 €. The worst performer is CPBDeg. For the total capacity fade $Q_{\textrm{loss}}$, CEmpDeg has the highest degradation in the summer and BNoDeg has the highest in the winter. The proposed ageing-aware CPBDeg achieves the lowest degradation in both seasons, maintaining a reasonable $C_{\textrm{grid}}$.
@@
In general, the planner that has the lowest $C_{\textrm{grid}}$ is the one with the highest $Q_{\textrm{loss}}$, but this is not always BNoDeg. The CEmpDeg fails to minimize the total capacity fade because of its model bias ($S^{M}_{a,t}\neq\tilde{S}^{M}_{a,t}$), in which it is rewarded to cycle at $\overline{SoC}_{b}$ and a calendar ageing independent of the $SoC_{b,t}$. Additionally, the linear BNoDeg planner fails to fulfill its predictions because of the high number of rejected actions during summer.
@@
Figure 10: Full equivalent cycles $FEC$ over time (left) and relative capacity fade $Q_{\textrm{loss}}$ over FECFEC (right).
@@
To dive deeper into the total storage usage, a quantitative analysis of the number of cycles done by the $b$ is necessary. Figure 10 presents the full equivalent cycles $FEC$ over time $t$, showing that the CPBDeg planners increase the EV usage $FEC_{\textrm{EV}}$ at the expense of doing fewer cycles with the BESS. When analyzing the capacity fade $Q_{\textrm{loss}}$ against the full eq. cycles $FEC$, Fig. 10, it is clear that the relative degradation per cycle ($\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$) of the CPBDeg is smaller in the winter than the used benchmarks. In summer, the BNoDeg has the lowest degradation per cycle due to its high number of rejects. The use of the EV s is the same for all planners, because the trajectory is mostly driven by user mobility. Finally, in both seasons, CEmpDeg presents the highest $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$. This appears to be a risky strategy due to a lack of consistency across seasons and objectives (minimizing degradation or minimizing grid costs). Lastly, even though the capacity fade is not significant in $T=1$ month, daily optimization can have a significant impact in the long term, as it was shown in [47, 46, 39]. As a final note, if the C-rate is increased ($\geq$1) and battery temperature $T$ is not constant, the degradation on a daily basis can be significantly higher.
@@
Finally, the computational time for the different strategies is presented in Fig. 11. Each sample is the total computational time it takes to solve Algorithm 2. Unexpectedly, BNoDeg has the lowest and most consistent $t_{\textrm{comp}}$ distribution, i.e., the smallest standard deviation. However, both CEmpDeg and CPBDeg planners have similar empirical distributions, maintaining overall fast computational time. Hence, the increase in modeling accuracy of PBROM is not prohibitively expensive when compared to its empirical counterpart. This is to be expected as the empirical ageing model is also non-linear and non-convex. In summary, this case study shows that:
  • 1.
    CPBDeg achieves the smallest $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ (best ageing control) and smallest total $Q_{\text{loss}}$.
  • 2.
    The computational cost of CPBDeg is not prohibitive when compared to its empirical benchmark CEmpDeg.
  • 3.
    BNoDeg achieves one of the smallest $C_{\text{grid}}$ with fast computational times, as expected.
5.2 Case Study II: Manageing different cathodes
To demonstrate the PB models’ flexibility and extended capabilities, the CPBDeg scheduler is tested using two similar battery packs of the same rated capacity $Q$ but using different cells. One is formed with NMC cells and the other with LFP. Since LFP cells have a lower rated capacity of $Q_{n}=2.3$ Ah and a lower $OCV$, the battery packs have more $N_{s/p}$ to have roughly the same pack-rated capacity as their NMC counterparts. The power limits $\overline{P}_{b,t},\ \underline{P}_{b,t}$ are also maintained to make an even comparison.
Figure 12: Cathode comparison simulation $SoC_{\textrm{EV}}$. Grey bands indicate driving periods.
The PB ageing models are suitable for both because they have graphite anodes [30, 15]. Nevertheless, they have different electrolytes. This is addressed by changing the electrolyte parameters in the model. The same model equations are used, but with different parameter values. This is a great advantage compared to the empirical fits presented in the literature. In the latter, the derived models are prone to overfitting to training conditions, delivering complex non-linear equations that can only be applied to specific chemistries and operating conditions.
@@ 5.2 Case Study II: Manageing different cathodes
The simulation results are presented in Figs. 12 - 14. Starting with the EV, Fig. 12, the operation is similar except for a few days in summer and winter in which the LFP decides to have deeper discharges than its NMC counterpart. Moving forward to the BESS, Fig. 13, in summer, the LFP scheduler is too ambitious, generating early storage depletion roughly 40% of the time, 10% more than the NMC cells. During the winter, total rejections are reduced to 25% and 10% for LFP and NMC, respectively.
Figure 14: Case Study 2 - LFP and NMC cells degradation analysis.
Table 2 presents the summary of performance for both cell types. Overall, the CPBDeg-LFP achieves lower total costs and ageing despite its high number of rejected actions. During winter $C_{\textrm{grid}}$ stays the same with 37% less degradation. However, in the summer, just by changing the cell type from NMC to LFP $C_{\textrm{grid}}$ is reduced by 25%, similar to BNoDeg in Section 5.1, and roughly 25% less capacity fade. This shows that CPBDeg rightly exploits its physical information of the system to achieve better performance.
@@
5.3 Case Study III: Manageing aged and fresh batteries
To demonstrate the flexibility and extended capabilities of the CPBDeg planner the scheduler is tested using two battery packs: the first is the fresh battery pack of NMC cells of Section 5.1 and the second is the same pack but with cells aged at $SoH=90\%$. Only one benchmark is used: a BNoDeg with no $SoH$ update. Thus, the BNoDeg EMS sees a perfectly healthy cell with rated capacity when in reality the battery pack is aged 10%. There is no empirical benchmark because the only parameter that can be update in it is the initial lifetime $t_{0,b}$, only affecting calendar ageing. The update is based on a 5% increase in series resistance $R_{b,0}$ and 10% decrease of the available Li content $z_{100\%,b}$ and its propagation with the equations of Section 3.3.2.
Figure 15 presents the degradation patterns for the 2 planners and different $SoH$. In the aged battery, the share of calendar ageing (against the total) within $i_{\text{SEI},b,t}$ is much smaller and thus the percentual $Q_{\textrm{loss}}$ is almost 35% smaller than in the new battery packs. In both seasons, the EV s patterns are similar for both BNoDeg and CPBDeg. The change in EV trajectories between aged and fresh cells shows a slight increase $FEC$, due to a smaller rated capacity, and a decrease in relative ageing, due to reduced calendar ageing. The impact on the BESS is more pronounced. In winter, CPBDeg has a smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ (upper-right hand side of the graph) for both aged and new cells, with BNoDeg doing more eq. cycles. In the winter CPBDeg still has a smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ in the new cell but not in the used cell, where this is achieved by the benchmark controller. In other words, as the battery ages, the predominance of calendar ageing within the SEI layer is diminished. Thus, the link between $i_{b,t}$ and $Q_{\text{loss},b,t}$ becomes stronger, due to the predominance of active material loss and the cyclic component of the SEI layer.
@@ 5.3 Case Study III: Manageing aged and fresh batteries
For the current simulation time $T$ of 1 month and average C-rate below 1C, the degradation slope may appear linear to the naked eye. However, in Fig. 15, the trajectory of the aged battery is also presented, and its average slope is smaller than the new battery degradation slope ($\frac{\Delta Q_{\textrm{loss},b}}{\Delta FEC_{b}}|_{SoH=90\%}<\frac{\Delta Q_{\textrm{loss},b}}{\Delta FEC_{b}}|_{SoH=100\%}$). This is due to the nonlinear dependency of $i_{\text{SEI},b,t}$ with $\sqrt{t}$. Thus, it is clear that even though each trajectory might appear linear, over longer simulation times of several months and years, the dependency is non-linear.
Finally, when adding the costs to the analysis, summarized in Table 3, it is noticed that CPBDeg achieves lower capacity fade $Q_{\textrm{loss}}$ than BNoDeg across all seasons and $SoH$. Even improving the total grid cost $C_{\textrm{grid}}$ as the cells degrade. In particular during summer, with a 35% observed improvement. On the other hand, BNoDeg worsens its performance as the cells degrade with higher $C_{\textrm{grid}}$, $Q_{\textrm{loss}}$ and model bias.
In summary, the performance of the proposed CPBDeg planner comparatively improves when using used cells, with respect to its linear BNoDeg benchmark. This is because the ageing PBROM s encapsulate the fade of calendar ageing ($\sqrt{t}$) and rise of cyclic SEI and AM ($i_{b,t},SoC_{b,t}$) as the battery ages.
@@
5.4 Limitations & Future works
@@
    Even though the terminal set is more flexible than the standard literature, the design is still arbitrary. It would be interesting to design the terminal conditions as part of a wider control scheme, such as seasonal optimization.
@@
    State observers are also necessary for all devices in an experimental application to correctly feed back the states $S_{a,t}$ to the optimization.
@@
    The thermal flow-based models used in this paper assume a fixed building temperature decided in an external system. This is the assumption behind $P_{\textrm{load}}^{\textrm{th}}$. Flexible thermal controls could also be integrated.
@@
    Uncertainty handling is limited since re-optimization frequency is only 24 hours. Future works include improving this through dynamic programming approaches and/or economic model predictive control.
@@ 6 Conclusions & Discussion
In summary, this paper presents an optimization-based day-ahead planner for residential multi-carrier energy systems that uses PBROM models to integrate battery ageing. Its integration with the thermal and mobility carriers is achieved by designing specific terminal sets for the individual storage systems, depending on their use and technology. The proposed CPBDeg planner can handle different cathode chemistries as well as batteries in different ageing states. The planner does this with lower degradation than the benchmarks. The PBROM integration comes at the expense of slightly higher computational times and grid cost in the case of NMC cells.
@@
In the first case study, it is shown how advanced PBROM can be used to reduce battery ageing while maintaining a grid cost comparable to benchmarks, in accordance with the literature for standalone utility-scale applications [41, 27, 65, 39, 38]. This is because the planner can better control the degradation per full equivalent cycle than the benchmarks.
@@
In the second case study, the proposed planner is equipped with battery packs of different cathode chemistries, and its performance is compared. Empirical models are cathode-specific and thus can’t handle multiple cell chemistries. The LFP battery pack has a lower total $Q_{\textrm{loss}}$ as per established knowledge and can achieve lower grid costs $C_{\textrm{grid}}$ than its NMC counterpart. Even more so, when considering that the relative cost $c_{\textrm{loss}}$ for both packs was the same, when in reality LFP packs have a lower cost than NMC packs. Thus, Section 5.2 is a conservative approximation, and LFP packs have the potential to enable even lower operation costs. The grid savings are achieved just by changing some simple model parameters taken from the literature. This flexibility is an essential feature because it allows the EMS to exploit any battery pack at hand fully.
The last case study showed how the proposed EMA handles aged and new batteries seamlessly, even improving performance as the battery ages, improving the summer grid cost by 35%. The reason is that the EMA can identify predominant degradation mechanisms from the ageing PBROM, exploiting the decrease of SEI importance over time and the rise of active material loss. Its linear counterpart, however, is unaware of the degradation, increasing the number of rejected actions. Summing up, grid costs can be reduced by using LFP battery packs or modifying objective weights, the latter at the expense of higher degradation. Finally, the integration of ageing PBROM s has a low impact on the HP and TESS, meaning that the almost 25% of grid cost savings of Section 5.2 are achieved without fundamentally modifying the thermal strategy. In applications where scalability and fast computation are key drivers, decomposition strategies could be explored to exploit this coupling.